OpenStudy (anonymous):

how to find the time required to fall an electron to proton when electron is attracted to proton?

OpenStudy (roadjester):

Dang, you're gonna need to specify that it's electromagnetism. I've only gotten to the "Mechanics" section of physics. :(

OpenStudy (jamesj):

What is the Coloumb force between the two particles? It is \[ F = \frac{ke^2}{r^2} \] where e is the charge on electron/proton, r is the distance between them and k is the Coloumb constant. Now if you want to solve the problem of how it would take for an electron and proton initially at rest to collide, you'll need to solve the (mechanics!) problem of for what time t is r(t) = 0. This is not an easy problem. You can find another working of that sort of problem here: http://openstudy.com/study?login#/updates/4f42fe64e4b065f388dc742a

OpenStudy (angelina):

see you have the force experienced by the electron due to proton which is F=k e^2/r^2 if you have the dist b/w them and they are at rest. now dividing it by mass of the electron you would get the accln and then u can calculate the time by the formula s=ut + 1/2at^2. but if they were not initially at rest then i dont know and if u come to know plz let me also know it

OpenStudy (jamesj):

(*correction: link I gave above is not correct; it is this one below) You can't naively apply s = ut + at^2/2 because here the acceleration is not constant precisely because F as a function of distance r is not a constant. See this link for how to solve the resulting differential equation: http://openstudy.com/study#/updates/4f0de661e4b084a815fcff56